# NL2KQL: From Natural Language to Kusto Query

### v2

@@ Abstract.
copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainccs: Information systems Language modelsccs: Computing methodologies Natural language generationccs: Human-centered computing Natural language interfaces
@@ 1. Introduction
Data is growing in volume and complexity, in turn, becoming harder to navigate even for the query language experts. Query languages are featureful and powerful; however they present a learning curve that may not be readily mastered by all users. Reducing technical obstacles would enable more data enthusiasts to access, understand and benefit from the data; hence, there is an ever increasing research on Natural Language (NL) Interface for Databases (NLIDB) to empower both technical and non-technical users to pose queries in NL and communicate with data. With Structured Query Language (SQL) being the most widely used query language, many NLQ-to-SQL solutions have been proposed over the years ranging from parsing-based approaches, to more recent Neural Machine Translation (NMT) techniques (Katsogiannis-Meimarakis and Koutrika, 2023). The two widely used NLQ-SQL datasets, WikiSQL (Zhong et al., 2017) and Spider (Yu et al., 2018b), and the recently released BIRD benchmark (Li et al., 2023), have led to nearly a thousand NLQ-to-SQL solutions.
Unlike SQL, which deals with structured data hence the name, Kusto Query Language (KQL) is designed for large semi-structured data such as logs, telemetry data, and time-series, which are commonly found in big data analytics platforms. KQL is designed to work with data lacking fixed or uniform structure with varying schemas within the same dataset and is shipped with a rich set of functions and operators for text and time-series making it highly effective to analyze complex patterns. These characteristics are blessings for KQL as a powerful query language and a curse in converting NL to KQL.
@@ 2.1. Retrieval-enhanced Generation
2.2. Few-shot learning
Few-shot learning, which focuses on models learning effectively from limited examples, has been an area of interest since 2004 and gaining momentum with the one-shot learning approach (Vinyals et al., 2017). The introduction of large pre-trained models has sparked a renewed focus, as seen in the famous Emergent Abilities work (Wei et al., 2022a). In-context Learning (ICL), a method introduced by OpenAI (Radford et al., 2019) that involves providing models with task descriptions in natural language, has further advanced the few-shot learning movement (Brown et al., 2020b).
@@ 2.3. NLQ-to-SQL
For a long while, rule-based methodologies and custom heuristics dominated the NLQ-to-SQL domain (Baik et al., 2019; Li and Jagadish, 2014; Saha et al., 2016; Sen et al., 2020). These methods typically employ established NLP techniques to deconstruct a question into a parsed tree. Based on the tree’s structure, specific rules are formulated to transform this tree into an SQL Abstract Syntax Tree (AST), which is subsequently converted into the executable SQL query. A critical aspect of this process is identifying and correctly assigning table and column names, as well as string literals, which involves a detailed examination of the database content to establish a correlation between the natural language elements and database schema components. Nonetheless, adapting these methods to new database domains often demands manual effort.
Recently, sketch-based learning techniques have introduced more generalizable solutions where a SQL sketch or template with placeholders is employed, further simplifying the task to merely filling in the details based on the natural language query. These methods bypasses the need for generating SQL syntax, focusing instead on the critical task of extracting and inserting relevant information into the predefined slots (Xu et al., 2017; Yu et al., 2018a).
In contrast, sequence-to-sequence models view SQL queries as a series of tokens, applying deep learning techniques. The primary challenge here is ensuring the syntactical correctness of the output SQL sequence. Notably, PICARD (Scholak et al., 2021) represents an advanced solution in this domain, employing a syntax-validation mechanism during the beam search process to exclude syntactically incorrect SQL fragments. Similarly, other research efforts focus on generating valid ASTs via a sequence-to-tree methodology, ensuring syntactical validity of the outputs, with RAT-SQL (Wang et al., 2020) being a prominent example. ValueNet discusses use of values along with schema information to improve the performance (Brunner and Stockinger, 2021). Katsogiannis et al. have composed a thorough review of NLQ-to-SQL research (Katsogiannis-Meimarakis and Koutrika, 2023).
@@ 3.1. Kusto Query Language (KQL)
Kusto Query Language (KQL) is a rich and powerful query language designed to analyze big data stored in Azure Data Explorer, a highly scalable data exploration service from Microsoft Azure [^1] https://learn.microsoft.com/en-us/azure/data-explorer/kusto/query/. KQL enables users to extract, manipulate, and analyze large volumes of data with high performance and ease of use. KQL is not just a query language but a tool for insights, offering a way to interact with vast datasets efficiently and intuitively. Unlike traditional SQL, KQL is optimized for unstructured and semi-structured data like logs, metrics, and events, which are common in modern applications and systems. Usecases of KQL are vast and diverse including Application Performance Monitoring (APM), Security Information and Event Management (SIEM), and Operation Intelligence.
@@
KQL is designed to be easy to read and write aiming at varying persona who use KQL in their day-to-day operations from Data Scientists, to DevOps and IT professionals, Security Analysts, Business Analysts, and Product Managers, some of whom have limited technical expertise. Consequently, a natural language textual interface that assists with writing KQL queries creates new opportunities for data analytics, helps further lower the barrier to entry for advanced operations and making KQL even more so accessible to a broader audience.
@@
<TableExpression>
| <operator> [parameters]
| <operator> [parameters]
// Data Retrieval
TableName
| take 100
// Filtering
TableName
| where ColName == Value
// Aggregation
TableName
| summarize Count = count() by ColName
// Sorting
TableName
| order by ColName asc
KQL, albeit simple, is a powerful data analytics tool with a wide range of aggregation functions  [^2] https://learn.microsoft.com/en-us/azure/data-explorer/kusto/query/aggregation-functions. The syntax is flexible and support various other syntactical forms and constructs that do not strictly adhere to the above mentioned pipeline pattern. These query forms are further discussed in Appendix B.
@@ 4.1. Semantic Data Catalog
A Semantic Data Catalog plays a pivotal role in encapsulating information about the structure, semantics, and contextual attributes of a database. The Schema Refiner and Query Refiner both leverage Data Catalog and the embeddings of its elements to make decisions for schema filtering and query repair (Figure 3).
Semantic Data Catalog provides annotations for tables, columns, type constraints, and, where applicable, enumerated values. The schema of the Data Catalog is as followed:
@@
The Semantic Data Catalog capture both structural and semantic information and is created via an automated process. For a given Kusto database, the schema structure, including table names, column names, and column types are retrieved, either directly from the database (.show database schema) or a third-party API. Once the structure is retrieved, available public documents are extracted from Azure Monitor Data Reference  [^3] https://learn.microsoft.com/en-us/azure/azure-monitor/reference/ including table descriptions, column descriptions, and in some cases, value descriptions (e.g. description of error codes). Generic data retrieval queries are executed on the Kusto database to obtain a sample of data for each table. This data is used as a proxy to extract unique values for the low-cardinality Enum columns. The acquired descriptions ensure that the Semantic Data Catalog contains both the structure and the semantics of the elements in database.
@@
The embeddings of the extended summaries of all elements are calculated via a similarity embedding model (Figure 3) and stored in a vector database for use at inference time. The extended summary is particularly helpful when dealing with columns like ”ErrorCode” which contain values numerical values (e.g. 16000 and 40008) necessitating contextual information for accurate understanding.
@@ 4.2. Schema Refiner
Overemphasis on Table Schema:
As highlighted in Section 3.3, the prevalent data redundancy and repetitive columns within tables present obstacles for the KQL generator. As the number of tables included in the schema increase, the likelihood of the autoregressive model incorrectly focusing on irrelevant tokens during generation rises.
@@
During inference, the schema is initially filtered based on the user’s permissions. Then, a maximum of $t$ similar tables are identified by calculating the cosine similarity between their embeddings and the NLQ embedding, where $t$ is a predefined hyper-parameter. All the columns in the selected tables are included in the schema. As discussed in Section 4.1, for some low-cardinality columns, which are denoted as Enum, the list of acceptable values are included in the prompt. However, to not pollute the context with irrelevant tokens, only the top $v_{\text{n}}$ values of a column whose embedding have close similarity with the NLQ embedding would be included in the schema.
@@ 4.3. Synthetic Few-shot Database
The NL2KQL solution uses a synthetic few-shot database (FSDB) to provide guidance for the LLM. The FSDB is synthesized for a target Kusto database and contains KQLs which are valid with respect to the database’s schema. The generated KQLs go through 3 rounds of validation to ensure sanity, syntactic and semantic correctness, and to filter out any ambiguous NLQs.
The synthetic KQL generation starts with random table sampling from the database schema. For 70% of queries only one table is sampled. For the remaining 30%, another random table is chosen to join with the first table. When choosing the joint table, in 50% of occurrences the table is only selected from the subset of tables which have similar column types. The final set of tables (i.e., one table without join, and two tables with join) are passed to the next step (Figure 2).
@@
An LLM is then prompted to generate a KQL query given the table set (schema) and the theme (Figure 2). The prompt contains instructions on Kusto syntax with query design options. The prompt also makes suggestions regarding the difficulty of the query and how to construct the final KQL including number of pipes.
Once the KQL query is generated, the same LLM is tasked to explain the KQL in imperative natural language to create an NLQ. The original KQL generation process is then repeated with an extended prompt which now includes the NLQ to generate a secondary KQL query. Both primary and secondary KQLs are validated with KQL parser for syntactic and semantic correctness and disregarded if they contain errors.
The initial and secondary KQLs are compared based on the Jaccard similarity of their tokens and rejected if their similarity falls below the 0.7 point. Accepted primary KQLs and their corresponding NLQs are added to the FSDB. Here, tokens are defined through whitespace splits. This validation ensures that the FSDB contains NLQ-KQL pairs for which NLQ is not ambiguous and is a true representation of its associated KQL query.
@@ 4.6. Query Refiner
To detect undefined identifiers, it evaluates if all identifiers exist within the accessible symbols at the given text position, aligning with the schema. When it encounters an undefined column name in a specified table, the function adjusts the identifier, substituting the incorrect segment with an appropriate column name from the same table, when such a column is available. To identify potential column replacements, it probes for column names within the same scope that share a prefix with the absent identifier. This method is similarly applied to resolve undefined variables in join operations, where it also considers the column names from the opposite side of the join as potential substitutes, assuming these are correctly defined. In extreme cases where multiple column identifiers are undefined, it tries to substitute the containing table with another table that encompasses all required columns.
When all of the above fail to fix the undefined identifier issue, Query Refiner uses embeddings of columns’ and tables’ descriptions, with a strict cosine similarity threshold of 0.9, to find substitute identifiers. As the undefined identifier lacks description, description of a column in the global scope which shares its name with the undefined column is used for embedding similarity matching.
@@
Figure 3. Overview of embedding stores. The NL2KQL solution uses three embeddings stores for tables, values, and few-shots. Embeddings are pre-calculated and used in inference to filter and select semantically relevant data catalog elements and few-shot examples.
@@ 5. EXPERIMENTS
Moreover, to create a more challenging environment with a reasonably large schema size, it was assumed that the user has access to all the tables in the Kusto database’s schema. During Schema Refinement, maximum number of tables, $t$, was set to 9. Maximum number of values per each Enum column, $v_{\text{n}}$, was set 5. The number of few-shots, $f$, was set to 5.
@@ 5.1. Evaluation Benchmark
Sentinel: The first schema originated from a Kusto database of Microsoft Sentinel, a Security Information and Event Management (SIEM) solution that delivers intelligent security analytics and threat intelligence across the enterprise. The Sentinel database has 23 tables. A team of security operations center analysts drafted 189 NLQ-KQL pairs targeted at the Sentinel database which are entirely used as the evaluation set.
Defender: The second schema is from Microsoft Defender XDR, an Extended Detection and Response (XDR) solution that coordinates detection, prevention, investigation, and response across endpoints, identities, email, and applications pre- and post-breach. The Defender database has 29 tables. Following the same process as the other database, 224 NLQ-KQL pairs were drafted as the Defender evaluation set.
@@
A summary of the schema, including list of tables and columns are available in Appendix D.
@@ 5.2.2. Online Metrics
The inherent data redundancy of logs and telemetries in Kusto databases and KQL’s flexibility in expressing the same query in different ways erodes uniqueness of ground-truth KQL that corresponds to a given NLQ. Online metrics, which measure the similarities of the query results, are intended to address the limitations of offline metrics, which only evaluate the syntactic and semantic similarity of the queries. The online metrics are as followed.
@@ 6. Results and Analysis
The proposed method is compared with the available baselines across the 8 mentioned metrics. Considering the many components of NL2KQL, a thorough ablation study is conducted where performance of alternatives to each component as well as the redacted versions of NL2KQL is evaluated. The results of offline and online metrics are summarized in Table 1 and Table 2, respectively. Across all metrics, we have found the $\text{Avg}_{\text{exec}}$ metric, which aggregates $\text{Row}_{\text{exec}}$ and $\text{Column}_{\text{exec}}$ to be the best indicator of performance.
@@
As highlighted in the results, all components of the NL2KQL considerably contribute to the end-to-end performance. Removing the schema (row 6), which originates from the Semantic Data Catalog, degrades $\text{Avg}_{\text{exec}}$ from 0.5800 down to 0.3888. Removing few-shots from this approach (row 6), yields the solution on row 7 with a 55% lower value of $\text{Avg}_{\text{exec}}$ at 0.1755, while adding the Schema Refiner to the mix (row 5) improves this metric by 40% to 0.5496. However, as the results demonstrate, in presence of the Schema Refiner (rows 1, 2, 3, 4), impact of few-shots saturate, yet, we still observe a 2% improvement from the 2-shot solution (row 3) to the main proposed solution (row 1) according to the $\text{Avg}_{\text{exec}}$ metric.
Results show that removing Query Refiner has a negative impact on both syntactic and semantic scores, reducing them to $0.9875$ and $0.9449$, respectively, which, in turn, negatively impacts the $\text{Avg}_{\text{exec}}$ metric as queries with errors fail to execute. Despite its positive impact, the Query Refiner made a mistake on two samples in the evaluation set slightly impacting the $\text{Filter}_{\text{col}}$ value of the NL2KQL method. Query Refiner is a deterministic post-processing component, therefore, we closely studied its behaviour on the evaluation set and observed that, on average it fixed 8.6 queries in the Defender database and 4.3 queries in the Sentinel database across the triplicate runs. Details of this study is reported in Appendix C.
For the Few-shot Selector component, two alternatives were explored: 2-shot (reducing few-shot examples from 5 to 2) and zero-shot (removing few-shot examples). We also conducted experiments with more few-shots up to 20 and we observed no further improvements to the performance which is in par with the reports in literature (Ahmed and Devanbu, 2023; Brown et al., 2020a). While NL2KQL consistently performed better across 7 out of the 8 metrics compared to its ablated versions, the zero-shot achieved better results on the $\text{Filter}_{\text{lit}}$ metric. We hypothesize that the few-shots could have influenced the literals’ format which negatively impacts this metric given our verbatim naive literal checks.
@@
In the ablation study of Schema Refiner, the entirety of the schema was included in the model’s context (row 5). In another experiment the schema was excluded from the prompt (row 6). Both scenarios demonstrated subpar performance compared to the proposed NL2KQL, further showcasing the effectiveness of Schema Refiner. We also experimented with higher values of $t$ (tables) which also negatively impacted the performance as it polluted model’s context with irrelevant information.
@@
Finally, the NLQ of the evaluation set were directly asked from the GPT-4 model, which is known to power the ChatGPT Plus service, which substantially under-performed compared to the NL2KQL solution, with $\text{Avg}_{\text{exec}}$ of 0.1077. Furthermore, to highlight the impact of the instructions in the Prompt Builder, the static sections of the NL2KQL prompt including KQL syntax guide and best practices were added to the prompt for GPT-4 which, on average, demonstrated $50\%$ better performance across all metrics.
@@
Table 1. Offline metrics. Comparison of the proposed method, NL2KQL, with baselines and ablated methods, highlighting the performance for each metric in bold.
@@
NL2KQL main 0.9933 0.9724 0.8338 0.6938 0.6051
NL2KQL w/o Query Refiner 0.9875 0.9449 0.8338 0.6943 0.6052
NL2KQL 2-shot 0.9866 0.9632 0.8303 0.6925 0.6000
NL2KQL Zero-shot 0.9783 0.9607 0.8372 0.6860 0.6122
NL2KQL w/o Schema Refiner 0.9875 0.9365 0.8233 0.6721 0.6101
NL2KQL w/o Schema 0.9816 0.6500 0.5886 0.5781 0.5679
GPT-4 with KQL Guide 0.9499 0.2523 0.3325 0.3900 0.5088
GPT-4 w/o KQL Guide 0.7494 0.1228 0.2561 0.3706 0.5649
Table 2. Online metrics. Comparison of the proposed method, NL2KQL, with baselines and ablated methods, highlighting the performance for each metric in bold.
@@
NL2KQL main 0.4441 0.7159 0.5800
NL2KQL w/o Query Refiner 0.4325 0.7137 0.5731
NL2KQL 2-shot 0.4297 0.6946 0.5622
NL2KQL Zero-shot 0.4391 0.6945 0.5668
NL2KQL w/o Schema Refiner 0.4169 0.6823 0.5496
NL2KQL w/o Schema 0.2822 0.4954 0.3888
GPT-4 with KQL Guide 0.1265 0.2245 0.1755
GPT-4 w/o KQL Guide 0.0823 0.1330 0.1077

### v3

@@ Abstract.
copyright: rightsretainedccs: Information systems Language modelsccs: Computing methodologies Natural language generationccs: Human-centered computing Natural language interfaces
@@ 1. Introduction
Query languages are feature-rich and powerful but often present a learning curve that not all users can easily master. Consequently, there is an ever increasing research on Natural Language (NL) Interface for Databases (NLIDB). With Structured Query Language (SQL) being the most widely used query language, many NLQ-to-SQL solutions have been proposed over the years ranging from parsing-based approaches, to more recent Neural Machine Translation (NMT) techniques (Katsogiannis-Meimarakis and Koutrika, 2023). The two widely used NLQ-SQL datasets, WikiSQL (Zhong et al., 2017) and Spider (Yu et al., 2018b), and the recently released BIRD benchmark (Li et al., 2023), have led to nearly a thousand NLQ-to-SQL solutions.
Unlike SQL, which deals with structured data hence the name, Kusto Query Language (KQL) is designed for large semi-structured data such as logs, telemetry data, and time-series, which are commonly found in big data analytics platforms. These characteristics are blessings for KQL as a powerful query language and a curse in converting NL to KQL.
@@ 2.2. NLQ-to-SQL
For a long while, rule-based methodologies and custom heuristics dominated the NLQ-to-SQL domain (Baik et al., 2019; Li and Jagadish, 2014; Saha et al., 2016; Sen et al., 2020). These methods typically employ established NLP techniques to deconstruct a question into a parsed tree. Based on the tree’s structure, specific rules are formulated to transform this tree into an SQL Abstract Syntax Tree (AST), which is subsequently converted into the executable SQL query. Nonetheless, adapting these methods to new database domains often demands manual effort.
Recently, sketch-based learning techniques have introduced more generalizable solutions where a SQL sketch or template with placeholders is employed. These methods bypasses the need for generating SQL syntax, focusing instead on extracting and inserting relevant information into the predefined slots (Xu et al., 2017; Yu et al., 2018a).
In contrast, sequence-to-sequence models view SQL queries as a series of tokens, applying deep learning techniques. The primary challenge here is ensuring the syntactical correctness of the output SQL sequence. Notably, PICARD (Scholak et al., 2021) represents an advanced solution in this domain, employing a syntax-validation mechanism during the beam search process ValueNet discusses use of values along with schema information to improve the performance (Brunner and Stockinger, 2021). Katsogiannis et al. have composed a thorough review of NLQ-to-SQL research (Katsogiannis-Meimarakis and Koutrika, 2023).
@@ 3.1. Kusto Query Language (KQL)
Kusto Query Language (KQL) is a powerful query language designed to analyze big data stored in Azure Data Explorer, a scalable data exploration service from Microsoft Azure [^1] https://learn.microsoft.com/en-us/azure/data-explorer/kusto/query/. KQL enables users to extract, manipulate, and analyze large datasets, making it more than just a query language but a tool for insights. Unlike traditional SQL, KQL is optimized for unstructured and semi-structured data like logs, metrics, and events, which are common in modern applications and systems.
@@
Despite its simplicity, KQL is a powerful data analytics tool with flexible syntax and various forms and constructs beyond the pattern. Further details on these query forms are discussed in Appendix B.
@@ 4.1. Semantic Data Catalog
A Semantic Data Catalog plays a pivotal role in encapsulating information about the structure, semantics, and contextual attributes of a database. The Schema Refiner and Query Refiner both leverage Semantic Data Catalog and the embeddings of its elements to make decisions for schema filtering and query repair (Figure 3).
Semantic Data Catalog provides annotations for tables, columns, type constraints, and, where applicable, enumerated values. The schema of the Semantic Data Catalog is as followed:
@@
The Semantic Data Catalog is created via an automated process. For a given Kusto database, the schema structure are retrieved, either directly from the database or a third-party API, and schema descriptions are extracted from Azure Monitor Data Reference  [^2] https://learn.microsoft.com/en-us/azure/azure-monitor/reference/ Generic data retrieval queries are executed on the Kusto database to obtain a sample of data for each table, allowing for the extraction of unique values for the low-cardinality Enum columns.
@@
The embeddings of the extended summaries of all elements are calculated via a similarity embedding model (Figure 3) and stored in a vector database for use at inference time.
@@ 4.2. Schema Refiner
During inference, the schema is initially filtered based on the user’s permissions. Then, a maximum of $t$ similar tables are identified by calculating the cosine similarity between their embeddings and the NLQ embedding, where $t$ is a predefined hyper-parameter. All the columns in the selected tables are included in the schema. As discussed in Section 4.1, for some low-cardinality columns, the list of acceptable values are included in the prompt. However, to not pollute the context with irrelevant tokens, only the top $v_{\text{n}}$ values of a column whose embedding have close similarity with the NLQ embedding would be included in the schema.
@@ 4.3. Synthetic Few-shot Database
The NL2KQL solution uses a synthetic few-shot database (FSDB) to guide the LLM. The FSDB is synthesized for a target Kusto database and contains KQLs which are valid with respect to the database’s schema. The generated KQLs go through 3 rounds of validation to ensure sanity, syntactic and semantic correctness, and to filter out any ambiguous NLQs.
The synthetic KQL generation starts with random table sampling from the database schema. For 70% of queries only one table is sampled. For the remaining 30%, a second table is chosen for a join, with half of these selected based on similar column types. The final set of tables (i.e., one table without join, and two tables with join) are passed to the next step (Figure 2).
@@
An LLM generate a KQL query given the table set (schema) and the theme (Figure 2). The prompt contains instructions on Kusto syntax with query design options.
Once the KQL query is generated, the same LLM is tasked to explain the KQL in imperative natural language (NLQ). This process is repeated with the NLQ included in the prompt to generate a secondary KQL query. Both primary and secondary KQLs are validated with KQL parser for syntactic and semantic correctness and disregarded if they contain errors.
The initial and secondary KQLs are compared using the Jaccard similarity of their tokens and rejected if their similarity falls below the 0.7 point. Accepted primary KQLs and their corresponding NLQs are added to the FSDB, ensuring that each NLQ accurately represents its KQL query without ambiguity.
@@ 4.6. Query Refiner
To detect undefined identifiers, Query Refiner uses embeddings of columns’ and tables’ descriptions, with a strict cosine similarity threshold of 0.9, to find substitute identifiers.
@@
Figure 3. Overview of embedding stores
@@ 5. EXPERIMENTS
Moreover, to create a more challenging environment with a reasonably large schema size, it was assumed that the user has access to all the tables in the Kusto database’s schema. During Schema Refinement, maximum number of tables, $t$, was set to 9. Maximum number of values per each Enum column, $v_{\text{n}}$, was set 5. The number of few-shots, $f$, was set to 2. The FSDB includes 16 few-shots for Defender and 19 for Sentinel.
@@ 5.1. Evaluation Benchmark
Sentinel: The first schema originated from a Kusto database of Microsoft Sentinel, a Security Information and Event Management (SIEM) solution that delivers intelligent security analytics and threat intelligence across the enterprise. The Sentinel database has 23 tables. A team of security operations center analysts drafted 197 NLQ-KQL pairs targeted at the Sentinel database which are entirely used as the evaluation set.
Defender: The second schema is from Microsoft Defender XDR, an Extended Detection and Response (XDR) solution that coordinates detection, prevention, investigation, and response across endpoints, identities, email, and applications pre- and post-breach. The Defender database has 29 tables. Following the same process as the other database, 230 NLQ-KQL pairs were drafted as the Defender evaluation set.
@@
The two evaluation sets for Defender and Sentinel, along with their respective Semantic Data Catalogs, are published and publicly accessible at https://github.com/¡to-be-released-post-acceptance¿. A summary of the schema, including list of tables and columns are available in Appendix C.
@@ 5.2.2. Online Metrics
The inherent data redundancy of logs and telemetries in Kusto databases and KQL’s flexibility in expressing the same query in different ways erodes uniqueness of ground-truth KQL that corresponds to a given NLQ. Online metrics, which measure the similarities of the query results, are intended to address the limitations of offline metrics. The online metrics are as followed.
@@ 6. Results and Analysis
The proposed method is compared with the available baselines across the 8 mentioned metrics. A thorough ablation study evaluate the performance of each component and their alternatives. The results of offline and online metrics are summarized in Table 1 and Table 2, respectively. Across all metrics, we have found the $\text{Avg}_{\text{exec}}$ metric, which aggregates $\text{Row}_{\text{exec}}$ and $\text{Column}_{\text{exec}}$ to be the best indicator of performance.
@@
As highlighted in the results, all components of the NL2KQL considerably contribute to the end-to-end performance. Specifically, removing the schema (row 6), which originates from the Semantic Data Catalog, degrades $\text{Avg}_{\text{exec}}$ from 0.6350 down to 0.4309. Furthermore, excluding the few-shots from this approach (row 6), yields the solution on row 7, which shows a 46% decrease in $\text{Avg}_{\text{exec}}$, bringing it to 0.2321. Conversely, adding the Schema Refiner to the mix (row 5) improves this metric by 8.7% from 0.5842.
Results show that removing Query Refiner has a negative impact on seven out of eight metrics, with no change in the Table metrics. This removal decreases the $\text{Avg}_{\text{exec}}$ by 1% to 0.6265.
For the Few-shot Selector component, two alternatives were explored: the all-shots approach (including all few-shot examples) and the zero-shot approach (removing few-shot examples). The results indicate that, in presence of the Schema Refiner (rows 1, 2, 3, 4), the impact of few-shots saturate. Nonetheless, we still observe a 1% improvement in the $\text{Avg}_{\text{exec}}$ metric from the all-shots solution (row 2) to the main proposed solution (row 1) with the addition of Few-shot Selector. However, this addition results in a slight decrease of 0.2% in both the Semantic and $\text{Column}_{\text{exec}}$ metrics. We hypothesize that while the Few-shot Selector enhances overall performance by providing targeted examples, it inadvertently introduces bias towards particular schema present in the few-shot examples. This bias can lead to a slight decrease in the model’s generalization. Despite NL2KQL consistently performed better across 7 out of the 8 metrics, the zero-shot achieved better results on the Table metric. We hypothesize that the few-shots could have influenced the table selection in a way that negatively impacts this metric.
@@
In the ablation study of Schema Refiner, the entirety of the schema was included in the model’s context (row 5). In another experiment the schema was excluded from the prompt (row 6). Both scenarios demonstrated subpar performance compared to the proposed NL2KQL, further showcasing the effectiveness of Schema Refiner.
@@
Finally, the NLQ of the evaluation set were directly asked from the GPT-4 model, which substantially under-performed compared to the NL2KQL solution, with $\text{Avg}_{\text{exec}}$ of 0.1150. Furthermore, to highlight the impact of the instructions in the Prompt Builder, the static sections of the NL2KQL prompt including KQL syntax guide and best practices were added to the prompt for GPT-4 which, on average, demonstrated $50\%$ better performance across all metrics.
@@
Table 1. Offline metrics. Comparison of the proposed method, NL2KQL, with baselines and ablated methods.
@@
NL2KQL main 0.9878 0.9603 0.8222 0.6992 0.6656
NL2KQL all-shots 0.9846 0.9627 0.8204 0.6965 0.6583
NL2KQL w/o Query Refiner 0.9862 0.9400 0.8222 0.6979 0.6647
NL2KQL Zero-shot 0.9830 0.9538 0.8414 0.6600 0.6304
NL2KQL w/o Schema Refiner 0.9838 0.9213 0.8040 0.6735 0.6262
NL2KQL w/o Schema 0.9854 0.6626 0.6765 0.5898 0.6334
GPT-4 with KQL Guide 0.9586 0.3406 0.3731 0.4159 0.5610
GPT-4 w/o KQL Guide 0.6513 0.1590 0.2238 0.3877 0.5639
Table 2. Online metrics. Comparison of the proposed method, NL2KQL, with baselines and ablated methods.
@@
NL2KQL main 0.5335 0.7365 0.6350
NL2KQL all-shots 0.5147 0.7394 0.6270
NL2KQL w/o Query Refiner 0.5270 0.7259 0.6265
NL2KQL Zero-shot 0.4989 0.6854 0.5922
NL2KQL w/o Schema Refiner 0.4795 0.6888 0.5842
NL2KQL w/o Schema 0.3331 0.5287 0.4309
GPT-4 with KQL Guide 0.1727 0.2914 0.2321
GPT-4 w/o KQL Guide 0.0803 0.1497 0.1150